# Accelerating Boolean Constraint Propagation for Efficient SAT-Solving on FPGAs

### v1

@@ V Experiments and Results
SAT Instance Davis et al [7] Thong et al [9] Our Design
@@
TABLE I: Comparison of BCP engine throughput (BCPs/s) with related work. Results reflect maximum theoretical throughput, achieved only data is fully available to BCP engines.
Figure 5: Breakdown of the total execution time across constituent components.
Related work calculates throughput (in BPCs performed per second), assuming full data availability: i.e., not taking into account software execution and communication/data transfer latency. Whilst this is a useful metric to assess hardware performance in isolation (and we report equivalent results in Table I), it does not accurately depict system performance; to do so, we break down the full execution in Figure 5 and evaluate speedup over vanilla software implementation, evaluating combinations of clause and variable sizes, with speedup depicted in Table II for meaningful combinations. For each combination, we also depict real throughput, in the form of BCPs/s averaged over total execution time (63 variables and 224 clauses is the theoretical upper bound, without the need for hot swapping). To evaluate the different effects of clause/variable sizes on execution, we fix one and vary the other, measuring total execution time: results are depicted in Figures 6 and 7.
@@
TABLE II: Varied clause/variable sizes and their impact on the relative speedup of hardware/software and the effective throughput of BCP engines.
Figure 6: Effect of increasing clauses size on total execution time, for 63 variables.
Figure 7: Effect of increasing variables size on total execution time, for 22400 clauses.
VI Conclusions
@@
Acknowledgments
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC)

### v2

@@ 5. Experiments and Results
SAT Instance Davis et al (Davis et al., 2008b) Thong et al (Thong and Nicolici, 2015) Our Design
@@
Table 1. Comparison of BCP engine throughput (BCPs/s) with related work. Results reflect maximum theoretical throughput, achieved only data is fully available to BCP engines.
Figure 4. Breakdown of the total execution time across constituent components.
Related work calculates throughput (in BPCs performed per second), assuming full data availability: i.e., not taking into account software execution and communication/data transfer latency. Whilst this is a useful metric to assess hardware performance in isolation (and we report equivalent results in Table 1), it does not accurately depict system performance; to do so, we break down the full execution in Figure 4 and evaluate speedup over vanilla software implementation, evaluating combinations of clause and variable sizes, with speedup depicted in Table 2 for meaningful combinations. For each combination, we also depict real throughput, in the form of BCPs/s averaged over total execution time (63 variables and 224 clauses is the theoretical upper bound, without the need for hot swapping). To evaluate the different effects of clause/variable sizes on execution, we fix one and vary the other, measuring total execution time: results are depicted in Figures 5 and 6.
@@
Table 2. Varied clause/variable sizes and their impact on the relative speedup of hardware/software and the effective throughput of BCP engines.
Figure 5. Effect of increasing clauses size on total execution time, for 63 variables.
Figure 6. Effect of increasing variables size on total execution time, for 22400 clauses.
6. Conclusions
@@
Acknowledgements.
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC).